OpenStudy (anonymous):

Jacqui has grades of 83 and 75 on her first two algebra tests. If she wants an average of at least 76, what possible scores can she make on her third test. A. greater than 70 B. 70 or greater c. 70 or less D less than 70

OpenStudy (skullpatrol):

Average = 76 = (83 + 75 + x)/3 228 = 83 + 75 +x 228 = 158 + x 70 = x this will give an average of 76 The question asks for "an average of at least 76,"